OpenStudy (anonymous):

The sum of three consecutive numbers is 84. What is the largest of these numbers? 26 27 28 29

OpenStudy (anonymous):

x+y+z=84 consecutively it basically saying 1+2+3=6, so knowing its asking for the largest number of the 3 variables, work downwards and guess and check to see what = 84.

OpenStudy (anonymous):

28?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

For example, A. 26+25+24=75 no the answer

OpenStudy (anonymous):

Try D.

OpenStudy (anonymous):

29+28+27=?